Past movie download sales for a company averaged 354 million a month, and varied according to a Normal distribution. The company wants to reduce the price to sell more downloads. After the price reduction, sales in the next 6 months averaged 398 million, with a standard deviation of 17. Write the hypothesis for test using alpha=0.01. Choose a test statistic. Calculate a test statistic. Find the p value.
Answer #1

given data are:-

sample mean () = 398

sample size (n) = 6

sample sd (s) = 17

hypothesis:-

test statistic be:-

degress of freedom = (n-1)= (6-1) = 5

p value be:-

[ in any blank cell of excel =T.DIST.RT(6.3399,5) press enter]

decision:-

p value = 0.0007 < 0.01 (alpha)

so, we reject the null hypothesis at 0.01 level of significance.

*** if you have any doubt regarding the problem please write it in the comment box.if you are satisfied please give me a LIKE if possible...